# Compatibility of Drinfeld presentations for split affine Kac–Moody quantum symmetric pairs

### v1

@@ 1.2. Reduced expressions for fundamental weights
In Step $1$ above, we compute explicit reduced expressions for the fundamental weights in the extended affine Weyl groups of classical types. Somewhat surprisingly, to our knowledge, such explicit expressions have not appeared in the literature before. As this result is interesting in its own right, we collect the expressions in a list below for quick reference, see Table 1.
@@ 1.3. Application to $q$-characters
In light of the aforementioned difficulties, we propose to define boundary $q$-characters directly via the Lu–Wang presentation instead. More precisely, we consider the generalized eigenspace decomposition of a finite dimensional representation with respect to the action of the operators $\Theta_{i,m}$, and let boundary $q$-characters encode the multiplicities of such eigenspaces. This is equivalent to taking the trace of a certain operator in the completion ${\mathbf{U}^{\imath}\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, yielding a map $\chi_{q}^{\imath}\colon\operatorname{Rep}\mathbf{U}^{\imath}\to{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$. We apply (1.1)–(1.2) to show that $\chi_{q}^{\imath}$ is compatible with the usual $q$-character map (Corollary 9.3). More precisely, we show that the diagram
@@
commutes if ${U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$ is endowed with an appropriate ‘twisting’ action of ${\mathbb{Z}}[Y_{i,a}^{\pm 1}]_{i\in I,a\in{\mathbb{C}}^{\times}}$. In particular, this result yields an easy way to compute the boundary $q$-characters of restriction representations.
@@ 2.1. Quantum affine algebras
We use standard notations and conventions regarding root systems, weight lattices, Weyl groups, etc., as in, e.g., [LW21, §3.1]. In particular, we let $\alpha_{i}$ ($i\in\mathbb{I}$) denote the simple roots of $\widehat{{\mathfrak{g}}}$; let $\theta$ denote the highest root, and $\delta$ the basic imaginary root; let $P$ and $Q$ denote the weight and root lattices of ${\mathfrak{g}}$, respectively; and let $\omega_{i}\in P$ ($i\in\mathbb{I}_{0}$) be the fundamental weights of ${\mathfrak{g}}$. Note that, according to Bourbaki’s conventions, $s_{\alpha_{i}}(\alpha_{j})=\alpha_{j}-a_{ji}\alpha_{i}$.
@@
We use the standard notation for divided powers, i.e., $(e^{\pm}_{i})^{(r)}=(e^{\pm}_{i})^{r}/[r]_{q_{i}}!$.
@@
The quantum loop algebra $U_{q}(L\mathfrak{g})$ is the quotient of $U_{q}(\widehat{\mathfrak{g}})$ by the ideal generated by the central element $K_{\delta}-1$, where $K_{\delta}$ is defined as in, e.g., [Bec94, §1.3]. Let us recall the “new” Drinfeld presentation of the quantum loop algebra $U_{q}(L\mathfrak{g})$. By [Dri87, Bec94], $U_{q}(L\mathfrak{g})$ is isomorphic to the algebra generated by $x^{\pm}_{i,k},h_{i,l},K^{\pm 1}_{i}$, where $k\in{\mathbb{Z}}$, $l\in{\mathbb{Z}}-\{0\}$ and $i\in\mathbb{I}_{0}$, subject to the following relations:
@@
It can be regarded as a version of the Drinfeld double of $U_{q}(\widehat{\mathfrak{n}}_{+})$. We consider $\widetilde{U}_{q}(L\mathfrak{g})$ as a $Q$-graded algebra with
@@ 2.4. The QSP braid group action
We have $T_{w}(B_{i})=B_{wi}$, $\mathbf{T}_{w}(B_{i})=B_{wi}$, for $i\in\mathbb{I}$ and $w\in\widetilde{W}$ such that $wi\in\mathbb{I}$.
@@ 2.5. Coideal structures
Given $\mathbf{c}=(c_{0},\cdots,c_{n})\in{\mathbb{C}}^{n+1}$, let $\mathbf{U}^{\imath}_{\mathbf{c}}$ be the quotient of $\widetilde{\mathbf{U}}^{\imath}$ by the two-sided ideal generated by $q_{i}^{-2}\mathbb{K}_{i}-c_{i}$. The induced map $\eta_{\mathbf{c},\mathbf{s}}\colon\mathbf{U}^{\imath}_{\mathbf{c}}\hookrightarrow\mathbf{U}$ gives $\mathbf{U}^{\imath}_{\mathbf{c}}$ the structure of a right coideal subalgebra of $\mathbf{U}$, with coproduct $\Delta_{\mathbf{c},\mathbf{s}}=\Delta\circ\eta_{\mathbf{c},\mathbf{s}}$. If $\mathbf{s}=(0,\cdots,0)$, we abbreviate $\eta=\eta_{\mathbf{s}}$ and $\eta_{\mathbf{c}}=\eta_{\mathbf{c},\mathbf{s}}$. Explicitly,
@@ 2.6. Rank one subalgebras
Let $\widetilde{\mathbf{U}}_{[i]}$ be the subalgebra of $\widetilde{\mathbf{U}}$ generated by $\mathbf{U}_{[i]}$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. The algebra isomorphism $\iota_{i}$ extends to an isomorphism $\iota_{i}\colon\widetilde{U}_{q}(L\mathfrak{sl}_{2})\to\widetilde{\mathbf{U}}_{[i]}$ sending $\mathbb{K}_{1}\mapsto K_{i}$ and $\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}$.
@@
We also note that $A_{-1}$ is sent to $\mathbf{T}_{\omega_{i}}(B_{i})$.
@@ 2.7. Factorization and coproduct in rank one
Let $\mathcal{H}$ denote the commutative subalgebra of $\widetilde{\mathbf{U}}$ generated by the coefficients of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$. We will refer to these coefficients as the Drinfeld–Cartan operators, and to $\mathcal{H}$ as the Drinfeld–Cartan subalgebra of $\widetilde{\mathbf{U}}$. Similarly, we refer to the coefficients of the series
@@
as $\imath$Drinfeld–Cartan operators, and to the commutative subalgebra $\mathcal{H}^{\imath}$ generated by them as the $\imath$Drinfeld–Cartan subalgebra of $\widetilde{\mathbf{U}}^{\imath}$. It is natural to ask how the two Drinfeld–Cartan subalgebras are related under the monomorphism (2.28). This question was answered in [Prz23] for quantum symmetric pairs of split affine type $\mathsf{A}$. Here we will need the rank one case of that result, i.e., when $\widetilde{\mathbf{U}}^{\imath}$ is isomorphic to the (universal) $q$-Onsager algebra.
@@ 3. Reduced expressions for fundamental weights
First, we need to introduce some notation. Let $w_{0}$ be the longest element of the finite Weyl group $W_{0}$ of ${\mathfrak{g}}$. Given $i\in\mathbb{I}_{0}$, let $w_{i}$ be the longest element of the Weyl group corresponding to the Dynkin diagram of ${\mathfrak{g}}$ with the node labelled by $i$ removed. Given $1\leq k<l$, let
@@
Proposition 3.1.
There is a bijection
$J\ \to\ \Lambda{-}\{1\},\qquad i\mapsto\omega_{i}w_{i}w_{0}.$
Proof.
See [Bou02, Ch. VI §2.3, Proposition 6] or [IM65, Proposition 1.18]. ∎
Proposition 3.2.
@@
where $\Delta^{+}$ are the positive roots of ${\mathfrak{g}}$.
Proof.
See [IM65, Proposition 1.23]. ∎
@@ 4.3. Type $\mathsf{C}_{n}^{(1)}$
The proof is similar to that of Lemma 4.3. ∎
@@ 5.1. Iterated $q$-brackets
(5.1) $P_{1}(y_{1})=$ $\ y_{1},\qquad$ $P_{k+1}(y_{1},\cdots,y_{k})=$ $\ P_{k}(y_{1},\cdots,y_{k-1},[y_{k},y_{k+1}]_{q}),$
(5.2) $P^{\prime}_{1}(y_{1})=$ $\ y_{1},\qquad$ $P^{\prime}_{k+1}(y_{1},\cdots,y_{k})=$ $\ [P^{\prime}_{k}(y_{1},\cdots,y_{k}),y_{k+1}]_{q}.$
@@ 5.2. Good polynomials
Let $P(B_{0},\cdots,B_{n})$ be any non-commutative polynomial in the variables $B_{0},\cdots,B_{n}$ with coefficients in ${\mathbb{C}}(\mathbb{K}_{i})_{i\in\mathbb{I}}$. Considering this polynomial as an element of $\widetilde{\mathbf{U}}$, and substituting for an occurence of a letter $B_{j}$ either $\widetilde{E}_{j}$ or $F_{j}$, we can write it as a sum of polynomials which are homogeneous in each of the variables $\widetilde{E}_{j}$ and $F_{j}$:
@@
From now on let us assume that each subterm satisfies $d_{0}^{+}+d_{0}^{-}=1$. Let $P_{+}$ be the sum of all subterms with $d_{0}^{+}=1$, and $P_{-}$ be the sum of all subterms with $d_{0}^{-}=1$. Moreover, let $P_{--}$ be the sum of all subterms which have maximal total degree and $d_{j}^{+}=0$ for all $j\in\mathbb{I}$.
@@ 7.3. Type $\mathsf{C}_{n}^{(1)}$
(7.21) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}1},\mathbb{P}(B_{n},P_{n{\shortminus}1}(B_{n{\shortminus}1},\cdots,B_{2},B_{1},B_{0})))),$
@@ 8.2. Explicit computation
$\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ \varepsilon^{\prime}_{k}\leq\varepsilon^{\prime}_{k+1}\end{subarray}}P_{i}(\tilde{e}_{i-1}^{\varepsilon^{\prime}_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon^{\prime}_{1}},P_{n{\shortminus}i}(F_{i+1},F_{i+2}\cdots,F_{n{\shortminus}1},P_{n{\shortminus}1}(\tilde{e}_{n}^{\varepsilon_{n}},\tilde{e}_{n-2}^{\varepsilon_{n-2}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}},F_{0}))).$
@@
which can be checked directly [^6] Such equalities can easily be verified using, e.g., GAP’s QuaGroup Package [dG03]. (they reduce to calculations in finite quantum groups of type $\mathsf{A}_{3}$ and $\mathsf{D}_{4}$, respectively). ∎

### v2

@@ 1.1. Proof strategy
We remark that the overall proof strategy is similar to that in [Prz23, §9]. However, each step of the proof is significantly harder than in type $\mathsf{A}$ case, mainly due to the higher complexity of reduced expressions for fundamental weights in other classical types. Another innovation is the concept of ‘good polynomials’, which allows us to systematically handle different types, and is also likely to find application in, e.g., the setting of quasi-split affine quantum symmetric pairs.
@@ 1.2. Reduced expressions for fundamental weights
In Step $1$ above, we compute explicit reduced expressions for the fundamental weights in the extended affine Weyl groups of classical types. Somewhat surprisingly, to our knowledge, such explicit expressions have not appeared in the literature before [^1] A recursive formula can be found in [Lus83].. As this result is interesting in its own right, we collect the expressions in a list below for quick reference, see Table 1.
@@ 1.3. Application to $q$-characters
In light of the aforementioned difficulties, we propose to define boundary $q$-characters directly via the Lu–Wang presentation instead. More precisely, we consider the generalized eigenspace decomposition of a finite dimensional representation with respect to the action of the operators $\Theta_{i,m}$, and let boundary $q$-characters encode the multiplicities of such eigenspaces. This is equivalent to taking the trace of a certain operator in the completion ${\mathbf{U}^{\imath}\otimes U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, yielding a map $\chi_{q}^{\imath}\colon\operatorname{Rep}\mathbf{U}^{\imath}\to{U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$, where $U_{q}(\widetilde{\mathfrak{h}})=\langle k_{i},h_{i,r}\mid i\in\mathbb{I}_{0},r\leq 0\rangle$ is a ‘half’ of the Drinfeld–Cartan subalgebra of $U_{q}(L{\mathfrak{g}})$. We apply (1.1)–(1.2) to show that $\chi_{q}^{\imath}$ is compatible with the usual $q$-character map (Corollary 9.3). More precisely, we show that the diagram
@@
commutes if ${U_{q}(\widetilde{\mathfrak{h}})}[\negthinspace[z]\negthinspace]$ is endowed with an appropriate ‘twisting’ action of ${\mathbb{Z}}[Y_{i,a}^{\pm 1}]_{i\in\mathbb{I}_{0},a\in{\mathbb{C}}^{\times}}$. In particular, this result yields an easy way to compute the boundary $q$-characters of restriction representations. For first results in the case of more general representations, we refer the reader to [LP25].
@@ 2.1. Quantum affine algebras
We use standard notations and conventions regarding root systems, weight lattices, Weyl groups, etc., as in, e.g., [LW21, §3.1]. In particular, we let $\alpha_{i}$ ($i\in\mathbb{I}$) denote the simple roots of $\widehat{{\mathfrak{g}}}$; let $\theta$ denote the highest root of ${\mathfrak{g}}$, and $\delta$ the basic imaginary root; let $P$ and $Q$ denote the weight and root lattices of ${\mathfrak{g}}$, respectively; and let $\omega_{i}\in P$ ($i\in\mathbb{I}_{0}$) be the fundamental weights of ${\mathfrak{g}}$. Note that, according to Bourbaki’s conventions, $s_{\alpha_{i}}(\alpha_{j})=\alpha_{j}-a_{ji}\alpha_{i}$.
@@
We use the standard notation for divided powers, i.e., $(e^{\pm}_{i})^{(r)}=(e^{\pm}_{i})^{r}/[r]_{q_{i}}!$. Given $\mu=\sum_{i\in\mathbb{I}}c_{i}\alpha_{i}\in{\mathbb{Z}}\mathbb{I}=\bigoplus_{i\in\mathbb{I}}{\mathbb{Z}}\alpha_{i}$, set
(2.6) $K_{\mu}=\textstyle\prod_{i\in\mathbb{I}}K_{i}^{c_{i}},\qquad K_{\delta}=K_{0}K_{\theta}.$
@@
The quantum loop algebra $U_{q}(L\mathfrak{g})$ is the quotient of $U_{q}(\widehat{\mathfrak{g}})$ by the ideal generated by the central element $K_{\delta}-1$. Let us recall the “new” Drinfeld presentation of the quantum loop algebra $U_{q}(L\mathfrak{g})$. By [Dri87, Bec94], $U_{q}(L\mathfrak{g})$ is isomorphic to the algebra generated by $x^{\pm}_{i,k},h_{i,l},K^{\pm 1}_{i}$, where $k\in{\mathbb{Z}}$, $l\in{\mathbb{Z}}-\{0\}$ and $i\in\mathbb{I}_{0}$, subject to the following relations:
@@
It can be regarded as a version of the Drinfeld double of $U_{q}(\widehat{\mathfrak{n}}_{+})$. Given $\mu\in{\mathbb{Z}}\mathbb{I}$, we define $\mathbb{K}_{\mu}$ in the same way as in (2.6). We consider $\widetilde{U}_{q}(L\mathfrak{g})$ as a $Q$-graded algebra with
@@ 2.4. The QSP braid group action
We have $T_{w}(e^{\pm}_{i})=e^{\pm}_{wi}$, $\mathbf{T}_{w}(B_{i})=B_{wi}$, for $i\in\mathbb{I}$ and $w\in\widetilde{W}$ such that $wi\in\mathbb{I}$.
@@ 2.5. Coideal structures
Given $\mathbf{c}=(c_{0},\cdots,c_{n})\in({\mathbb{C}}^{\times})^{n+1}$, let $\mathbf{U}^{\imath}_{\mathbf{c}}$ be the quotient of $\widetilde{\mathbf{U}}^{\imath}$ by the two-sided ideal generated by $q_{i}^{-2}\mathbb{K}_{i}-c_{i}$. The induced map $\eta_{\mathbf{c},\mathbf{s}}\colon\mathbf{U}^{\imath}_{\mathbf{c}}\hookrightarrow\mathbf{U}$ gives $\mathbf{U}^{\imath}_{\mathbf{c}}$ the structure of a right coideal subalgebra of $\mathbf{U}$, with coproduct $\Delta_{\mathbf{c},\mathbf{s}}=\Delta\circ\eta_{\mathbf{c},\mathbf{s}}$. If $\mathbf{s}=(0,\cdots,0)$, we abbreviate $\eta=\eta_{\mathbf{s}}$ and $\eta_{\mathbf{c}}=\eta_{\mathbf{c},\mathbf{s}}$. Explicitly,
@@ 2.6. Rank one subalgebras
Let $\widetilde{\mathbf{U}}_{[i]}$ be the subalgebra of $\widetilde{\mathbf{U}}$ generated by $\mathbf{U}_{[i]}$, $\mathbb{K}_{i}^{\pm 1}$ and $(\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1})^{\pm 1}$. The algebra isomorphism $\iota_{i}$ extends to an isomorphism $\iota_{i}\colon\widetilde{U}_{q}(L\mathfrak{sl}_{2})\to\widetilde{\mathbf{U}}_{[i]}$ sending $\mathbb{K}_{1}\mapsto\mathbb{K}_{i}$ and $\mathbb{K}_{0}\mapsto\mathbb{K}_{\delta}\mathbb{K}_{i}^{-1}$.
@@
We also note that $A_{1,-1}$ is sent to $\mathbf{T}_{\omega_{i}}(B_{i})$.
@@ 2.7. Factorization and coproduct in rank one
Let $\mathcal{H}$ denote the commutative subalgebra of $\widetilde{\mathbf{U}}$ generated by the coefficients $\phi^{\pm}_{i,\pm r}$ of the series $\boldsymbol{\phi}^{\pm}_{i}(z)$. We will refer to these coefficients as the Drinfeld–Cartan operators, and to $\mathcal{H}$ as the Drinfeld–Cartan subalgebra of $\widetilde{\mathbf{U}}$. Similarly, we refer to the coefficients of the series
@@
as $\imath$Drinfeld–Cartan operators [^2] One has $\boldsymbol{\grave{\Theta}}_{i}(z)=(q_{i}-q_{i}^{-1})\acute{\Theta}_{i}(z)$, where $\acute{\Theta}_{i}(z)=\sum_{r\geq 0}\acute{\Theta}_{i,r}z^{r}$ is the series from [LW21, (2.10)], whose coefficients are the imaginary root vectors from [BK20], rescaled by $-q_{i}^{-2}$., and to the commutative subalgebra $\mathcal{H}^{\imath}$ generated by them as the $\imath$Drinfeld–Cartan subalgebra of $\widetilde{\mathbf{U}}^{\imath}$. It is natural to ask how the two Drinfeld–Cartan subalgebras are related under the monomorphism (2.29). This question was answered in [Prz23] for quantum symmetric pairs of split affine type $\mathsf{A}$. Here we will need the rank one case of that result, i.e., when $\widetilde{\mathbf{U}}^{\imath}$ is isomorphic to the (universal) $q$-Onsager algebra.
@@ 3. Reduced expressions for fundamental weights
First, we need to introduce some notations. Let $w_{0}$ be the longest element of the finite Weyl group $W_{0}$ of ${\mathfrak{g}}$. Given $i\in\mathbb{I}_{0}$, let $w_{i}$ be the longest element of the Weyl group corresponding to the Dynkin diagram of ${\mathfrak{g}}$ with the node labelled by $i$ removed. Given $1\leq k<l$, let
@@
Proposition 3.1 ([Bou02, Ch. VI §2.3, Proposition 6], [IM65, Proposition 1.18]).
There is a bijection $J\longrightarrow\Lambda{-}\{1\}$ sending $i\mapsto\omega_{i}w_{i}w_{0}$.
Proposition 3.2 ([IM65, Proposition 1.23]).
@@
where $\Delta^{+}$ is the set of the positive roots of ${\mathfrak{g}}$.
@@ 4.3. Type $\mathsf{C}_{n}^{(1)}$
The lemma follows by a direct calculation. ∎
@@ 5.1. Iterated $q$-brackets
(5.1) $P_{1}(y_{1})=$ $\ y_{1},\qquad$ $P_{k+1}(y_{1},\cdots,y_{k+1})=$ $\ P_{k}(y_{1},\cdots,y_{k-1},[y_{k},y_{k+1}]_{q}),$
(5.2) $P^{\prime}_{1}(y_{1})=$ $\ y_{1},\qquad$ $P^{\prime}_{k+1}(y_{1},\cdots,y_{k+1})=$ $\ [P^{\prime}_{k}(y_{1},\cdots,y_{k}),y_{k+1}]_{q}.$
@@ 5.2. Good polynomials
Let $P(B_{0},\cdots,B_{n})$ be any non-commutative polynomial in the variables $B_{0},\cdots,B_{n}$ with coefficients in ${\mathbb{C}}(\mathbb{K}_{i})_{i\in\mathbb{I}}$. Considering this polynomial as an element of $\widetilde{\mathbf{U}}$, and substituting for an occurrence of a letter $B_{j}$ either $\widetilde{E}_{j}$ or $F_{j}$, we can write it as a sum of polynomials which are homogeneous555 For example, suppose that $P(B_{0},\cdots,B_{n})=B_{2}^{2}B_{1}$. Then $B_{2}^{2}B_{1}=\widetilde{E}^{2}_{2}\widetilde{E}_{1}+\widetilde{E}_{2}^{2}F_{1}+(\widetilde{E}_{2}F_{2}\widetilde{E}_{1}+F_{2}\widetilde{E}_{2}\widetilde{E}_{1})+(\widetilde{E}_{2}F_{2}F_{1}+F_{2}\widetilde{E}_{2}F_{1})+F_{2}^{2}\widetilde{E}_{1}+F_{2}^{2}F_{1},$ i.e., the monomial $B_{2}^{2}B_{1}$ becomes a sum of six homogeneous polynomials. in each of the variables $\widetilde{E}_{j}$ and $F_{j}$:
@@
We will now introduce some technical assumptions and definitions with view to proving Proposition 5.5. From this point on let us assume that each subterm satisfies $d_{0}^{+}+d_{0}^{-}=1$. Let $P_{+}$ be the sum of all subterms with $d_{0}^{+}=1$, and $P_{-}$ be the sum of all subterms with $d_{0}^{-}=1$. Moreover, let $P_{--}$ be the sum of all subterms which have maximal total degree and $d_{j}^{+}=0$ for all $j\in\mathbb{I}$.
@@ 7.3. Type $\mathsf{C}_{n}^{(1)}$
(7.21) $\mathbf{T}_{\omega^{\prime}_{i}}(B_{i})=P_{i}(B_{i{\shortminus}1},\cdots,B_{1},P_{n{\shortminus}i}(B_{i+1},\cdots,B_{n{\shortminus}1},\mathbb{P}(B_{n},P_{n}(B_{n{\shortminus}1},\cdots,B_{2},B_{1},B_{0})))),$
@@ 8.2. Explicit computation
$\sum_{\begin{subarray}{c}\varepsilon_{k}\leq\varepsilon_{k+1},\\ \varepsilon^{\prime}_{k}\leq\varepsilon^{\prime}_{k+1}\end{subarray}}P_{i}(\tilde{e}_{i-1}^{\varepsilon^{\prime}_{i-1}},\cdots,\tilde{e}_{1}^{\varepsilon^{\prime}_{1}},P_{n{\shortminus}i}(F_{i+1},F_{i+2},\cdots,F_{n{\shortminus}1},P_{n{\shortminus}1}(\tilde{e}_{n}^{\varepsilon_{n}},\tilde{e}_{n-2}^{\varepsilon_{n-2}},\cdots,\tilde{e}_{2}^{\varepsilon_{2}},F_{0}))).$
@@
which can be checked directly [^9] Such equalities can easily be verified using, e.g., GAP’s QuaGroup Package [dG03]. For an explicit code, see https://github.com/lijr07/identities_in_quantum_groups. (they reduce to calculations in finite quantum groups of type $\mathsf{A}_{3}$ and $\mathsf{D}_{4}$, respectively). ∎